# ML4H 2025 Template: Proceedings Track

### v3

@@ Abstract
Clinical workflows are fragmented as a patchwork of scripts and task-specific networks that often handle triage, task selection, and model deployment. These pipelines are rarely streamlined for data science pipeline, reducing efficiency and raising operational costs. Workflows also lack data-driven model identification (from imaging/tabular inputs) and standardized delivery of model outputs. In response, we present a practical, healthcare-first framework that uses a single vision–language model (VLM) in two complementary roles. First (Solution 1), the VLM acts as an aware model-card matcher that routes an incoming image to the appropriate specialist model via a three-stage workflow (modality $\rightarrow$ primary abnormality $\rightarrow$ model-card id). Checks are provided by (i) stagewise prompts that allow early exit via None/Normal/Other and (ii) a stagewise answer selector that arbitrates between the top-2 candidates at each stage, reducing the chance of an incorrect selection and aligning the workflow with clinical risk tolerance. Second (Solution 2), we fine-tune the VLM on specialty-specific datasets ensuring a single model covers multiple downstream tasks within each specialty, maintaining performance while simplifying deployment. Across gastroenterology, hematology, ophthalmology, and pathology, our single-model deployment matches or approaches specialized baselines.
Compared with pipelines composed of many task-specific agents, this approach shows that one VLM can both decide and do. It may reduce effort by data scientists, shorten monitoring, increase the transparency of model selection (with per-stage justifications), and lower integration overhead.
@@
Figure 1: Three-stage model-card matching workflow (modality $\rightarrow$ abnormality $\rightarrow$ model card id).
@@ 1 Introduction
Machine learning (ML) models in healthcare face significant operational friction when moving from research prototypes to clinical practice. An estimated 90% of medical AI models never reach clinical deployment Poddar et al. (2024), and production-ready models can take weeks to months to integrate, with over half of organizations reporting 8–90 days to deploy a single model Mage AI (2025). These delays stem from several bottlenecks, and we provide two key examples here. First, to select an appropriate model design (architecture, training objective, etc.) for a given dataset, the data scientist must identify the attributes that should guide the design. In tabular data these attributes are explicit; in clinical images they are latent in the pixels—such as modality (e.g., CT) and clinical indication (e.g., acute intracranial hemorrhage)—and must first be extracted, standardized, and quality-checked before they can inform design choices. As a result, streamlined model selection for clinical imaging is often overlooked and warrants dedicated attention Willemink et al. (2020). Second, each chosen model is an operational artifact: health systems integrate, validate, monitor, and recalibrate AI solutions one by one, so developing and maintaining task-specific models multiplies the cost burden and slows adoption Brady et al. (2024).
We address these operational issues with two linked solutions: the first assists data scientists in selecting the appropriate model, and the second outlines a practical first step to shorten time-to-deployment once that selection is finalized.
@@
Early medical vision–language efforts aligned images and text to enable cross-modal understanding and transfer. Recently, MedGemma Sellergren et al. (2025) and UMIT Yu et al. (2025) showed that medical VLMs can approach specialized systems while maintaining generality, motivating workflows that both interpret the input and decide what to do next. As such, we adopt a medical VLM as an aware, auditable selector that reads an input image and chooses an appropriate model card via a three-stage pipeline (modality $\rightarrow$ primary abnormality $\rightarrow$ model-card id). At each stage, the VLM proposes top candidate models, and an answer selector applies a calibrated threshold that can abstain and, when appropriate, terminate early to prevent error cascades. Decisions at each stage (VLM output, any abstention, and routing to the next stage) are visible to users and logged in a consistent format, yielding transparent selection tied to model-card semantics Mitchell et al. (2019) and safe selective prediction Geifman and El-Yaniv (2019). This process accelerates development: We believe that choosing the appropriate model for a given clinical image hinges on two key pieces of information—imaging modality and disease type—details that are not always apparent to data scientists.
@@
Prior work shows that domain-pretrained medical encoders, when fine-tuned on target tasks, can match strong baselines across multiple datasets Zhang et al. (2022); Mei et al. (2022). In healthcare, specialty-level foundation models have also demonstrated broad coverage within a specialty—outperforming narrow, dataset-specific models across many tasks (e.g., pathology’s Prov-GigaPath achieves SOTA on 25/26 tasks across 31 tissue types) Xu et al. (2024). Operationally, fewer, broader specialty models reduce the number of artifacts a health system must validate, secure, and monitor, lowering deployment burden compared with one-model-per-dataset estates Brady et al. (2024). In line with past work, we fine-tune the same calibrated VLM at the specialty level—gastroenterology, hematology, ophthalmology, and pathology—rather than per dataset. Our goal is to match task-specific model performance while keeping everything on a single, easy-to-maintain deployment stack. Once a model card is finalized by the joint decision of the data scientist and the Solution 1 selection workflow, the data scientist can adapt the VLM to the downstream use case rather than designing a new model. This adaptation consists of tailoring the VLM prompt to the use case and eliminates the time otherwise spent building and integrating a new architecture.
Taken together, these solutions link selection and deployment in a single, calibrated workflow—a VLM that can route and execute. Compared with recent multi-agent pipelines (Shimgekar et al., 2025b, a), this minimalist design may reduce data-science workload and reduce human error. As an example, AI-driven triage has been shown to cut chest-radiograph reporting delays from 11.2 to 2.7 days in simulation while maintaining diagnostic performance Annarumma et al. (2019). Similarly, our approach may shorten the path from triage to clinical use and increase throughput without sacrificing performance.
@@ 2 Related Work
Matching models based on an appropriate dataset has been explored at multiple layers of care delivery, such as patient-facing systems, specialty care, etc. Patient-facing systems (e.g., symptom checkers like Buoy Health  buoy health (2025)) narrow likely conditions and suggest next steps. Platforms such as Garner Health  Garner (2025) match patients to specialists using outcomes data. Similarly, LLM-driven controllers (HuggingGPT  Shen et al. (2023)) show how a coordinator can parse requests, consult model descriptions, and delegate to expert models. This pattern maps naturally to healthcare model cards as standardized summaries of intended use and performance Mitchell et al. (2019). Clinical imaging “orchestration” seeks to run the appropriate algorithm on the selected scan Aidoc (2024). However, prior work does not provide calibrated, stage-wise routing that is explicitly grounded in model-card semantics, with visible abstention and early termination and per-stage decision logs for audit. In the first solution, we address this gap with a three-stage router that makes every decision—VLM output, abstention, and route—user-visible and logged, cutting routing errors while preserving safety via careful prompt design and stage-specific selection logic.

### v4

# One VLM, Two Roles: Stage-Wise Routing and Specialty-Level Deployment for Clinical Workflows

@@ Abstract
Clinical ML workflows are often fragmented and inefficient: triage, task selection, and model deployment are handled by a patchwork of task-specific networks. These pipelines are rarely aligned with data-science practice, reducing efficiency and increasing operational cost. They also lack data-driven model identification (from imaging/tabular inputs) and standardized delivery of model outputs. We present a framework that employs a single vision–language model (VLM) in two complementary, modular roles. First (Solution 1): the VLM acts as an aware model-card matcher that routes an incoming image to the appropriate specialist model via a three-stage workflow (modality $\rightarrow$ primary abnormality $\rightarrow$ model-card ID). Reliability is improved by (i) stage-wise prompts enabling early termination via None/Other and (ii) a calibrated top-2 answer selector with a stage-wise cutoff. This raises routing accuracy by +9 and +11 percentage points on the training and held-out splits, respectively, compared with a baseline router, and improves held-out calibration (lower ECE). Second (Solution 2): we fine-tune the same VLM on specialty-specific datasets so that one model per specialty covers multiple downstream tasks, simplifying deployment while maintaining performance. Across gastroenterology, hematology, ophthalmology, pathology, and radiology, this single-model deployment matches or approaches specialized baselines.
Together, these solutions reduce data-science effort through more accurate selection, simplify monitoring and maintenance by consolidating task-specific models, and increase transparency via per-stage justifications and calibrated thresholds. Each solution stands alone, and in combination they offer a practical, modular path from triage to deployment.
@@
Figure 1: The first stage in the router workflow (modality detection).
Figure 2: The second stage in the router workflow (disease identification).
Figure 3: The third (final) stage in the router workflow (model-card matching).
@@ 1 Introduction
Machine learning (ML) models in healthcare face significant operational friction when moving from research prototypes to clinical practice. An estimated 90% of medical AI models never reach clinical deployment Poddar et al. (2024), and production-ready models can take weeks to months to integrate, with over half of organizations reporting 8–90 days to deploy a single model Mage AI (2025). These delays stem from several bottlenecks, and we provide two key examples here. First, to select an appropriate model design (architecture, training objective, etc.) for a given dataset, the data scientist must identify the attributes that guide model design. In tabular data these attributes are explicit; in clinical images they are latent in the pixels—such as modality (e.g., CT) and clinical indication (e.g., acute intracranial hemorrhage)—and must first be extracted, standardized, and quality-checked before they can inform design choices. As a result, streamlined model selection for clinical imaging is often overlooked and warrants dedicated attention Willemink et al. (2020). Second, health systems tend to integrate, validate, monitor, and recalibrate AI solutions one by one, so developing and maintaining task-specific models multiplies the cost burden and slows adoption Brady et al. (2024).
We address these operational issues with two linked—but independently evaluated—solutions: the first assists data scientists in selecting the appropriate model, and the second shortens time-to-deployment by unifying task-specific AI models into broader, specialty-level models that support multiple tasks. While Solution 1 could inform Solution 2 in future systems, our implementation treats them separately; Solution 2 does not consume outputs from Solution 1.
@@
Early medical vision–language efforts aligned images and text to enable cross-modal understanding and transfer. Recently, MedGemma Sellergren et al. (2025) and UMIT Yu et al. (2025) demonstrated that medical VLMs can approach specialized systems while maintaining generality, motivating workflows that both interpret the input and decide what to do next. Accordingly, we adopt a medical VLM as an aware, auditable selector that reads an input image and chooses an appropriate model card via a three-stage pipeline (modality $\rightarrow$ primary abnormality $\rightarrow$ model-card ID). At each stage, the VLM proposes top candidates, and an answer selector applies a calibrated threshold that either selects the most likely candidate (our baseline) or, if warranted, the second most likely candidate. Decisions at each stage (VLM output, any abstention, and routing to the next stage) are visible to users and logged in a consistent format, yielding transparent selection tied to model-card semantics Mitchell et al. (2019) and safe selective prediction Geifman and El-Yaniv (2019). This process accelerates development: we posit that choosing the appropriate model for a given clinical image hinges on two key pieces of information—imaging modality and disease type—details that are not always apparent to data scientists.
@@
Prior work shows that domain-pretrained medical encoders, when fine-tuned on target tasks, can match strong baselines across multiple datasets Zhang et al. (2022); Mei et al. (2022). In healthcare, specialty-level foundation models have also demonstrated broad coverage within a specialty—outperforming narrow, dataset-specific models across many tasks (e.g., pathology’s Prov-GigaPath achieves SOTA on 25/26 tasks across 31 tissue types) Xu et al. (2024). Operationally, fewer, broader specialty models reduce the number of artifacts a health system must validate, secure, and monitor, lowering deployment burden compared with one-model-per-dataset estates Brady et al. (2024). In line with past work, we fine-tune the same calibrated VLM at the specialty level—gastroenterology, hematology, ophthalmology, pathology, and radiology—rather than per dataset. Our goal is to match task-specific model performance while keeping everything on a single, easy-to-maintain deployment stack. To adapt the VLM to multiple AI tasks, one must include task-specific prompts in the VLM’s training data (e.g., a task-specific user prompt for colonoscopy lesion classification and a task-specific user prompt for polyp bounding-box detection) to build a specialty model that supports both classification and bounding-box (BBox) prediction. Importantly, in our experiments, Solution 2 is trained and evaluated independently of Solution 1.
Taken together, these solutions connect selection and deployment in a calibrated workflow—a VLM that can route and execute. Compared with recent multi-agent pipelines (Shimgekar et al., 2025b, a), this minimalist design may reduce data-science workload and human error. As an example, AI-driven triage has been shown to cut chest-radiograph reporting delays from 11.2 to 2.7 days in simulation while maintaining diagnostic performance Annarumma et al. (2019). Similarly, our approach may shorten the path from triage to clinical use and increase throughput without sacrificing performance.
@@ 2 Related Work
Selecting the appropriate model based on a dataset has been explored at multiple layers of care delivery, such as patient-facing systems, specialty care, etc. Patient-facing systems (e.g., symptom checkers like Buoy Health  buoy health (2025)) narrow likely conditions and suggest next steps. Platforms such as Garner Health  Garner (2025) match patients to specialists using outcomes data. Similarly, LLM-driven controllers (HuggingGPT  Shen et al. (2023)) show how a coordinator can parse requests, consult model descriptions, and delegate to expert models. This pattern maps naturally to healthcare model cards as standardized summaries of intended use and performance Mitchell et al. (2019). Clinical imaging “orchestration” seeks to run the appropriate algorithm on the selected scan Aidoc (2024). However, prior work does not provide calibrated, stage-wise routing that is explicitly grounded in model-card semantics, with visible abstention and early termination and per-stage decision logs for audit. In the first solution, we address this gap with a three-stage router that makes every decision—VLM output, abstention, and route—user-visible and logged, cutting routing errors while preserving safety via careful prompt design and stage-specific selection logic.
@@
3 Solution 1: Model–Card Matching with Early Termination
To integrate the medical VLM as an automated model selector, we employ a three-stage prompting pipeline. At each stage, the model receives a stage-specific instruction and context and returns a textual answer. The stages progressively narrow the information needed to select the appropriate model card. After each response, an answer-selection step ranks candidate answers by likelihood, retains the top two, and chooses between them; for Stage 1 only, it may also terminate early. Figures 1, 2, and 3 illustrate the overall workflow in order; Algorithm 1 provides pseudocode for the answer-selection strategy; and Algorithm 2 explains how we derive stage-wise cutoffs that parameterize Algorithm 1. Below, we first outline the prerequisites and then describe the selection process for each stage.
Prerequisite 1: Backbone VLM choice
In Section 1 (Solution 1) we described a generic medical VLM; here, we adopt MedGemma. MedGemma was pretrained on a multimodal medical mixture spanning radiology (CXR/CT/MRI), ophthalmology, dermatology, and histopathology [^1] Example datasets: MIMIC-CXR 231,483 images; EyePACS 199,258; internal dermatology 51,049; PAD-UFES-20 2,047; internal histopathology 32.55M patches; CT-US1 59,979 slices; MRI-US1 47,622 Sellergren et al. (2025). and reports competitive benchmarks relative to prior work Sellergren et al. (2025).
Prerequisite 2: Model-card repository
To match a clinical image to the correct model, we maintain a diverse repository of model cards spanning multiple modalities (e.g., CT, MRI) and AI tasks (e.g., classification, grading). This repository currently includes 84 model cards over 26 unique modalities: brain CT scan, brain MRI scan, breast histopathology (H&E) scan, breast mammography scan, breast ultrasound scan, cervix cytology scan, cervix histology scan, cervix histopathology (H&E) scan, chest CT scan, colonoscopy scan, colon histopathology (H&E) scan, esophageal gastroscopy scan, fundus photography scan, gallbladder ultrasound scan, kidney CT scan, kidney histopathology scan, kidney MRI scan, kidney ultrasound scan, liver CT scan, liver ultrasound scan, lung histopathology (H&E) scan, prostate histopathology (H&E) scan, skin dermatology scan, skin dermoscopy scan, skin histopathology (H&E) scan, and a skin multimodal panel (dermoscopy+RCM+H&E). Rather than storing task-specific models, each candidate is indexed by a string key—the model-card ID (a.k.a. model code; e.g., MODEL_04, which might refer to an ophthalmology model)—and by two minimal descriptors that make it discoverable by the router: (i) a short task caption describing what the model does (e.g., “segments tumor regions in histopathology images”), and (ii) the modality, i.e., the scan type on which the model was trained (e.g., “breast histopathology scan”).
Stage 1: Modality identification (Figure 1).
The goal of the first stage is to determine the imaging modality or scan type of the input image. We prompt MedGemma with a system message that casts it as a medical-imaging specialist and supply a list of admissible modalities drawn from the model-card repository. Concretely, we iterate over all model cards, extract each card’s modality field, deduplicate the values to form a unique set, and inject that set into the prompt. The prompt instructs: (1) if the image is not a medical scan, respond with None; (2) if it is a medical scan, choose the single most appropriate scan name from the list that matches the image content; and (3) if none of the listed modalities perfectly fits, respond with Other. This prompting strategy explicitly conditions the model to use the tokens None or Other when a confident match is absent. For example, given an endoscopic image of the colon, the correct output might be “colon colonoscopy scan”; for a non-medical photograph (e.g., a picture of a cat), the model should output None. Early termination occurs only at this stage: if the selected token is None or Other, the pipeline halts.
Stage 2: Primary finding detection (Figure 2).
In the second stage, the model examines the image (unaware of the router’s result for Stage 1) and attempts to identify the primary abnormality or finding, if any. We prompt MedGemma with an instruction such as: “Identify the single most likely disease or finding apparent in this scan. If no disease is present, respond only with: ‘Normal.’” This stage is essentially an image-based diagnosis step for a high-level finding. For example, given a colonoscopy image, the model might answer “Polyp” if one is seen; given a chest X-ray, it might answer “Normal” if nothing abnormal stands out, or a broad condition like “Pneumonia” if detected. The prompting ensures that the special token “Normal” is used to indicate the absence of any visible pathology.
Stage 3: Model-card selection (Figure 3).
In the final stage, the model selects an appropriate card from the repository. We iterate over all cards, filter to those whose modality matches the Stage 1 choice, and pass the resulting list of (model_card_id, task_caption) pairs to the prompt. For example, if Stage 1 predicts breast histopathology scan, the candidate set might be: [MODEL_01: Classifies benign vs. malignant findings in breast histopathology slides., MODEL_04: ].
To provide sufficient context for disambiguation, we include the Stage 1 modality and the Stage 2 primary abnormality as structured fields in the prompt (e.g., Patient scan details: breast histopathology scan; Primary abnormality: carcinoma). To make this a pure selection task, we mask the input image so that MedGemma acts as a language model solving a multiple-choice problem over the candidate list. The prompt instructs the model to: (i) review each candidate’s task caption; (ii) select a model only if its task explicitly matches both the identified modality and abnormality; (iii) reply None if no option fits; and (iv) when a match exists, return the model’s code (card ID) plus a one-line justification of the fit.
Cutoff Tuning and Answer Selection.
Even with careful prompting (e.g., return the best-matched, restrictive cues like only, and explicit abstentions such as None), the router can still pick suboptimal first tokens or miss safe abstentions. Empirically, we observe that the second-most likely first token can sometimes yield a more faithful full answer. We therefore arbitrate between the top two candidates using a simple, stage-specific cutoff on the runner-up probability.
Selector logic.
At each stage, the model forms a distribution over the first token. We take the top two tokens and decode full answers conditioned on each, producing $\{(a_{i},p_{i})\}_{i=1}^{2}$ with $p_{1}\!\geq p_{2}$. This pair is passed to a transparent selector that prefers the runner-up only when its probability is sufficiently large. The rule returns both the chosen answer $\hat{a}$ and the index $y\in\{\textsc{first},\textsc{second}\}$. When applying Algorithm 1 in Stage 1 only, if the selected token is literally None or Other, we halt and return None as the router’s final output (early termination to avoid extra compute).
[…]
2
Given $\mathcal{C}=\{(a_{i},p_{i})\}_{i=1}^{|V|}$ with $p_{1}\!\geq\!p_{2}$, cutoff $\tau$.
Output $\hat{a}$ (selected answer), $y\in\{\textsc{first},\textsc{second}\}$.
3
  1. 1.
    Set $\hat{a}\!\leftarrow a_{1}$ and $y\!\leftarrow\textsc{first}$.
  2. 2.
    If $p_{2}\!\geq\!\tau$ and $a_{2}\!\neq\!a_{1}$ then
    1. set $\hat{a}\!\leftarrow a_{2}$ and $y\!\leftarrow\textsc{second}$ (prefer runner-up when $p_{2}$ is large).
-2pt
Algorithm 1 Answer selector (per stage)
Learning the cutoff in a supervised fashion.
We gathered 102 diverse (image, per-stage top-2 predictions, label (First/Second)) pairs over the training set and 18 pairs over the held-out set. All samples were collected from Wikimedia Commons and share a Creative Commons Attribution–ShareAlike (CC BY-SA) license. For each stage $k\!\in\!\{1,2,3\}$, the training split provides the correct index (first vs. second). We learn a single cutoff $\tau_{k}$ via a one-pass sweep over the sorted runner-up probabilities $p_{2}$, which returns all optimal intervals and a canonical midpoint for deployment. Applying Algorithm 2 with the $p_{2}$-based selector (Alg. 1) yields the optimal intervals and representative cutoffs in Table 1. Intervals are open on the right; wide plateaus indicate robustness.
[…]
2
Given Training rows for stage $k$ with runner-up
probabilities $p_{2}$ and labels $y\!\in\!\{\textsc{first},\textsc{second}\}$.
Output All optimal cutoff intervals for $\tau_{k}$; canonical $\tau_{k}^{\star}$
(midpoint of the widest interval).
3
  1. 1.
    Partition events: $A\!\leftarrow\!\{p_{2}:\,y=\textsc{second}\}$,  $B\!\leftarrow\!\{p_{2}:\,y=\textsc{first}\}$.
  2. 2.
    Sort once: merge $A\cup B$ ascending, consuming equal values together (event list).
  3. 3.
    Initialize score: place $\tau$ just left of 0; $S_{0}\!\leftarrow\!|A|$.
  4. 4.
[…]
    1. (a)
      Crossing an $A$-value: a previously correct flip-to-second becomes incorrect $\Rightarrow S\!\leftarrow\!S-1$.
    2. (b)
      Crossing a $B$-value: a previously incorrect stay-with-first becomes correct $\Rightarrow S\!\leftarrow\!S+1$.
  5. 5.
    Track maxima: record $S_{\max}$ and every open interval between successive events where $S=S_{\max}$.
  6. 6.
    Return: these intervals are optimal; set $\tau_{k}^{\star}$ to the midpoint of the widest optimal interval.
-2pt
Algorithm 2 Cutoff sweep (training split, stage $k$)
Table 1: Optimal cutoff intervals and representative $\tau_{k}^{\star}$ learned on the training split.
Stage Optimal interval(s) $\tau_{k}^{\star}$
Stage 1 — Modality [0.417100, 0.434750) 0.425925
Stage 2 — Disease [0.436800, 0.437400) 0.437100
Stage 3 — Final [0.001100, 0.001334); [0.001985, 0.003030) 0.002508
Per-stage accuracy (training and held-out).
We compare a simple baseline (always choose first) with the cutoff router (Alg. 1 using $\tau_{k}^{\star}$ from Alg. 2). The $p_{2}$-based selector is transparent and stage-local. Cutoffs learned by a single sweep improve final-stage correctness on both splits (approximately +10 pp on training split; approximately +6 pp on held-out split), with modest gains at earlier stages. The inference-time Stage 1 abstention guard reduces unnecessary computation without affecting cutoff learning or the evaluations above.
Table 2: Per-stage accuracy with and without cutoff.
Split Stage / Task Baseline acc Cutoff acc (ours)
Training Stage 1 — Modality 0.9216 0.9412
Stage 2 — Disease 0.7647 0.7745
Stage 3 — Final 0.4608 0.5588
Held-out Stage 1 — Modality 0.8889 0.8889
Stage 2 — Disease 0.6667 0.6667
Stage 3 — Final 0.5556 0.6111
3.1 Router Results: Cutoff vs. Baseline
Setup. We compare the router’s final decision stage with and without a cutoff on the tuning splits (training: 102, held-out: 18). For each image, the router emits a predicted label (either a MODEL_# or None) together with a predicted confidence in $[0,1]$, i.e., the probability the router assigns to the emitted label at decision time. Accuracy is measured by exact match between the predicted label and the ground-truth label. Calibration is summarized with the Expected Calibration Error (ECE) using 10 equal-width, right-closed bins:
$\mathrm{ECE}\;=\;\sum_{b=1}^{10}\frac{n_{b}}{N}\,\bigl|\mathrm{acc}_{b}-\mathrm{conf}_{b}\bigr|\quad\text{(lower is better).}$
Table 3: Final decision stage results. Accuracy is exact match; ECE uses 10 equal-width, right-closed bins.
Split / Method Accuracy ECE
Training — Baseline (no cutoff) 0.392 0.415
Training — Cutoff (ours) 0.480 0.490
Held-out — Baseline (no cutoff) 0.444 0.390
Held-out — Cutoff (ours) 0.556 0.372
Reliability curves. Figures 4 and 5 plot accuracy vs. confidence per bin for the cutoff router and the baseline, along with the ideal y=xy{=}x diagonal. On the held-out split, the cutoff curve lies closer to the diagonal (lower ECE). On the training split, the cutoff router improves accuracy but is mildly under-calibrated at mid/high confidence, increasing ECE. Overall, Cutoff improves end-task correctness on both splits (+8.8 pp on training split; +11.1 pp on held-out split) and achieves better held-out calibration (ECE $0.372$ vs. $0.390$). A practical advantage, visible in Figures 4 and 5, is that when confidence exceeds $0.5$, accuracy under Cutoff tends to increase monotonically with confidence; in contrast, the baseline exhibits mixed behavior, where confidence does not reliably track accuracy. Furthermore, on the held-out split the cutoff curve closely mirrors the shape observed on the training split—aside from a slight deviation in the first (lowest-confidence) range—indicating that the reliability profile measured on the training data largely transfers to unseen data. In practice, this suggests that development-time reliability estimates are predictive of deployment-time reliability.
Figure 4: Reliability (Accuracy vs. Confidence) on training split. Dashed line: ideal y=xy{=}x. The cutoff router improves accuracy but is slightly under-calibrated in mid/high-confidence bins.
Figure 5: Reliability (Accuracy vs. Confidence) on held-out. The cutoff router improves both accuracy and calibration (lower ECE) compared to the baseline.
4 Solution 2: Specialty-Specific Model for Downstream Deployment
Building on the rationale in Section 1 (Solution 2), we fine-tune the same backbone VLM adopted in Section 3MedGemma—across five specialties (gastroenterology, hematology, ophthalmology, pathology, and radiology) to simplify maintenance and deployment rather than pursuing state-of-the-art performance on each task. Importantly, Solution 2 is implemented and evaluated independently; it does not consume outputs from Solution 1, though the two are conceptually complementary.
For each specialty we compare: (a) an external task-specific benchmark (typically the published state-of-the-art for that dataset), (b) the base MedGemma in zero-shot mode (no domain adaptation), and (c) MedGemma after specialty-specific fine-tuning (denoted “S–S” in the table). This tests whether a single specialty-tuned MedGemma can substitute for the external benchmark while retaining a single deployment stack. If successful, a fixed VLM plus targeted prompt engineering and supervised fine-tuning may reduce the need to design, validate, and operate separate architectures—saving compute and shortening triage-to-deployment cycles Annarumma et al. (2019). All results are summarized in Table 4.
Datasets
For gastroenterology, we use the dataset introduced by Li et al. (2021), an annotated compilation of three public sources: CVC-ColonDB Bernal et al. (2012), GLRC Mesejo et al. (2016), and KUMC (80 colonoscopy video sequences from the University of Kansas Medical Center). For hematology, we use BloodMNIST Acevedo et al. (2020), organized into eight blood cell categories. For ophthalmology, we employ OCTMNIST Kermany et al. (2018) (four retinal OCT classes) and RetinaMNIST Dataset (2020) (five retinopathy grades). For pathology, we use TissueMNIST Ljosa et al. (2012) (eight kidney cortex cell types), PathMNIST Kather et al. (2019) (nine colorectal tissue types), and a breast cancer histopathology set following Cruz-Roa et al. (2014). Lastly, for radiology, we use BreastMNIST Al-Dhabyani et al. (2020) (binary classification of breast images: normal vs. malignant), DeepLesion Yan et al. (2017) (eight different body organs), Organ(A-B-C)MNIST Bilic et al. (2023); Xu et al. (2019) (11 body organs in axial, coronal, and sagittal views, respectively), PneumoniaMNIST Kermany et al. (2018); Kermany (2018) (pneumonia vs. normal chest images), and ChestMNIST Wang et al. (2017) (14 disease labels in X-ray images). Train/validation/test sizes are summarized in Table 4.
Fine-tuning results
For each specialty-specific case (five in total), we use a parameter-efficient fine-tuning (PEFT) method, QLoRA Dettmers et al. (2023). As shown in Table 4, specialty-specific fine-tuning substantially narrows the gap to specialist benchmarks. While the base MedGemma performs modestly in zero-shot mode, targeted specialty adaptation yields competitive results across modalities and objectives. This supports the premise that one VLM can act as a generalist engine per specialty, reducing the need for separate per-task architectures and thereby saving compute and shortening triage-to-deployment cycles. We also provide public example inference results for each specialty-specific fine-tuning case here.
5 Discussion
We presented two linked clinical solutions: (1) a single, calibrated medical VLM that runs an aware model-card–matching workflow, routing an input image to a model-card ID—or abstaining and terminating when appropriate—and (2) the same VLM used as a specialty-specific downstream model with performance competitive with task-specific models. The solutions are complementary but decoupled in our implementation: Solution 2 neither requires nor consumes Solution 1 outputs. For Solution 1, the workflow is designed to streamline selection for data scientists by providing: safety and trust (stage-wise decisions are visible throughout the pipeline), robustness (fewer incorrect selections via the combined use of stage-wise prompts and answer-selector logic), scalability (flexibility to add stages, descriptors, or model cards, or to swap in a more mature selector), and speed (a minimalist design with fewer moving parts than recent multi-agent pipelines (Shimgekar et al., 2025b, a)). For Solution 2, specialty-specific models offer a practical first step toward automated deployment. We make this step concrete by shifting effort from designing new architectures to engineering task-specific prompts that align the VLM’s outputs with downstream objectives. In practice, this process can increase throughput for data scientists and clinicians without sacrificing performance.
Limitations and future work. MedGemma is pretrained on a fixed set of modalities and tasks; accordingly, it may not generalize to imaging types or findings outside its training distribution. In short, its inference and selection decisions are tied to its training data: if a modality or pathology is unseen or underrepresented, the model may fail to recognize it and either choose an imperfect model or default to “Other/None.” This risk can be mitigated by enlarging the pretraining corpus or by adopting continual-learning methods that add new domains while preserving prior knowledge (e.g., Li and Hoiem (2017)).
While our system outputs justifications for model selection, the interpretability of the VLM’s internal reasoning remains limited (a common issue with large neural models). Techniques such as chain-of-thought prompting or explicability constraints could be applied so the model provides brief, stage-wise reasoning—further improving transparency (see Wei et al. (2022)). Another design choice concerns the selection of thresholds $\tau_{1},\tau_{2},\tau_{3}$. We derive the optimal cutoff thresholds using Algorithm 2, which maximizes the per-stage score given (i) our stage-specific prompting strategy; (ii) the diversity and structure of the model-card repository defined in Section 3; and (iii) the workflow’s preference for the true final output—either None or a model-card ID—across different input images. In practice, these thresholds may require adjustment across clinical contexts. For high-accuracy applications using the same modality and disease labels as the training set, a workflow designer might prefer None over the closest model-card match, resulting in different labels for Stage 3 tuning.
Finally, our pipeline has three stages because repository lookup is primarily based on modality and a single primary finding. As we incorporate more complex model cards (e.g., some models might handle combinations of findings or require patient metadata), the prompt design may need to extend to additional stages or a more free-form query. Fortunately, the answer-selector concept generalizes to any number of stages by examining the confidence of alternative responses. We outline two complementary directions for deployment: (1) automate fine-tuning end to end—so deployment proceeds with minimal data-scientist intervention—by automatically selecting the dataset linked to the model card chosen in Solution 1 and automatically generating the prompts required for fine-tuning; and (2) reduce inference latency by applying pruning methods tailored to large language models (e.g., SparseGPT Frantar and Alistarh (2023)).